OpenStudy (anonymous):

I'm needing help (refresher) on power series and series in general. Calc II or Differential Equations, either way. First problem: Find all values of x for which this series converges: ∑(4^n)(x^n)(n+1)/(n+9); n0=1

OpenStudy (anonymous):

I'd rather if you said how to do it, rather than the answer.

OpenStudy (agreene):

\[\large \sum_{n=0}^{\infty}\frac{(4^n)(n+1)(x^n)}{(n+9)}\] is that correct?

OpenStudy (anonymous):

yeah, exactly right.

OpenStudy (anonymous):

sorry, i should have posted the pic

OpenStudy (anonymous):

Unfortunately I don't have my calc textbook, or it wouldn't be a problem. :(

OpenStudy (anonymous):

The ratio test would take care of this i believe. Let:\[a_n=\frac{4^n\cdot x^n(n+1)}{n+9}\] and you want to figure out when:\[\lim_{n\rightarrow \infty}\left|\frac{a_{n+1}}{a_n}\right|<1\]

OpenStudy (anonymous):

ok, so it would be for values of x that make that true.

OpenStudy (anonymous):

right.

OpenStudy (agreene):

I hate series sooo much >.<

OpenStudy (anonymous):

Basically we end up with the power series converging if:\[|4x|<1 \Longrightarrow |x|<\frac{1}{4}\]

OpenStudy (anonymous):

lol! I actually got that far before, just didn't know where to go from there.

OpenStudy (anonymous):

my differential equations book doesn't cover series very well

OpenStudy (anonymous):

and yeah, I find series a pain, too.

OpenStudy (anonymous):

you gotta go back to a Calculus book rather than a Diff Eqn book. Thats what i had to do lol.

OpenStudy (anonymous):

let's see if The Calculus Lifesaver says enough… :D

OpenStudy (anonymous):

hm… "it" (online homework software) doesn't like [-1/4,1/4] as the interval of convergence

OpenStudy (anonymous):

try ( , ) instead of brackets. The brackets give the impression that its ok if x = 1/4, but thats not the case.

OpenStudy (anonymous):

it doesn't take that either. (current answer is (-.25,.25)

OpenStudy (anonymous):

i cant figure out why it wouldnt take that...

OpenStudy (anonymous):

by the nth term test, I get (4^n)*(x^n)

OpenStudy (anonymous):

me neither. Might this be a case of needing to apply multiple tests?